OpenStudy (ladylauren):

Evaluate [(51 + 3) − 3^2] ÷ 9 ⋅ 2. 52 13.3 2.5 10

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (sunnnystrong):

Recall that Order of Operations --- PEMDAS [(51 + 3) − 3^2] ÷ 9 ⋅ 2. [(54)-9]/9*2 45/18 --> =?

OpenStudy (ladylauren):

2.5

OpenStudy (skullpatrol):

45/9 first please

OpenStudy (ladylauren):

5

OpenStudy (skullpatrol):

Now multiply by 2

OpenStudy (ladylauren):

10

OpenStudy (skullpatrol):

Yes

OpenStudy (skullpatrol):

Do all multiplications and divisions inorder from left to right

OpenStudy (ladylauren):

51+3=54 54-9=45 45/9=5 5*2=10 so the answer is 10

OpenStudy (skullpatrol):

You got! Good job :-)

OpenStudy (sunnnystrong):

@skullpatrol ... But you would multiply 9 & 2 first?

OpenStudy (skullpatrol):

That^ would be working from right to left

OpenStudy (sunnnystrong):

oh wait just kidding haha @skullpatrol XD

OpenStudy (sunnnystrong):

Cool